oteedee:

Thrifty supply store normally sells candy bars for $0.59. Today they are running a special 2 candy bars for $0.99. How much would you save if you bought 10 candy bars at the sale price rather than at regular price?

jhonyy9:

any idea ?

jhonyy9:

like a first step you need calcule ten candy bars how much cost for the regular price 2. step you need calcule ten candy bars how much cost for this sale price rater 3. step subtract from the regular price cost of ten candy bars the sale price rater cost of ten candy bars hope you understand this way

jhonyy9:

ok. ?

jhonyy9:

1. how much will cost ten candy bars for $0,59 per one

jhonyy9:

than 1 candy bar cost $0,59 so ten candy bars will cost 10*0,59 = ?

jhonyy9:

please collaborate

jhonyy9:

courage i like help you

jhonyy9:

so 10*0,59 = 5,9 what mean that 10 candy bars for the regular price will cost $5,9 now you need calcule the price of 10 candy bars for this sale price so this will be 5*0,99 = 4,95 what mean that 10 candy bars for this sale price will cost $4,95 how much would you save ? 5,9 - 4,95 = 0,95 what mean that will save $0,95 hope helped